I need help with B (FUTA and SUTA are incorrect).BMX Company has one employee. FICA Social Security taxes are 6.2% of the first $127,200 paid to its employee, and FICA Medica
Answer #1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 800.00 6.20% 49.60
FICA-Medicare 800.00 1.45% 11.60
FUTA 600.00 3.10% 18.60
SUTA 600.00 2.90% 17.40
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2100.00 6.20% 130.20
FICA-Medicare 2100.00 1.45% 30.45
FUTA 2100.00 3.10% 65.10
SUTA 2100.00 2.90% 60.90
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 6300.00 6.20% 390.60
FICA-Medicare 8000.00 1.45% 116.00
FUTA 0.00 3.10% 0.00
SUTA 0.00 2.90% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-6400 = $600
Note: Earnings taxable for Social Security in (c) = 127200-120900 = $6300